OpenStudy (julezzzzz):

Combine like terms. What is a simpler form of the expression? -3(-4y + 3) + 7y A. 19y - 9 B. 10y C. -19y + 3 D. -19y - 9

Nnesha (nnesha):

first distribute parentheses by -3 \[\large\rm \color{red}{a}(b+c)= \color{red}{a} \cdot b + \color{red}{a} \cdot c =\color{Red}{a}b +\color{Red}{a}c\]

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\)\(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

Nnesha (nnesha):

ty ty